Find the area bounded by the curve 𝑦 = (π‘₯^2)/4

and π‘₯ + 4 = 2y


1
Expert's answer
2021-06-21T16:32:55-0400
"y=\dfrac{x^2}{4}"

"x+4=2y=>y=\dfrac{1}{2}x+2"

"\dfrac{x^2}{4}=\dfrac{1}{2}x+2"

"x^2-2x-8=0"

"(x+2)(x-4)=0"

"x_1=-2, x_2=4"

"A=\displaystyle\int_{-2}^{4}(\dfrac{1}{2}x+2-\dfrac{x^2}{4})dx="

"=\big[\dfrac{x^2}{4}+2x-\dfrac{x^3}{12}\big]\begin{matrix}
 4 \\
 -2
\end{matrix}"

"=4+8-\dfrac{16}{3}-(1-4+\dfrac{2}{3})=9(units^2)"

"A=9" square units